OpenStudy (anonymous):

log(subscript4)2x=5 solve for x! help!

OpenStudy (amistre64):

log4(2x) = 5 right?

OpenStudy (anonymous):

mhm but the 4 is a subscrpt

OpenStudy (amistre64):

the 4 is the base... i know :)

OpenStudy (anonymous):

\[\log_ab = c \iff a^c = b \]

OpenStudy (anonymous):

yes ha how i sovle is

OpenStudy (amistre64):

log base"4" is can be written in shorthand log4

OpenStudy (anonymous):

k so 2x=4^5

OpenStudy (amistre64):

you can put a base 4 under each side right?

OpenStudy (amistre64):

good, just like that lol

OpenStudy (anonymous):

then divided by 2?

OpenStudy (amistre64):

now what is 4*4*4*4*4?

OpenStudy (anonymous):

ohhh uhh 1024

OpenStudy (amistre64):

is it? lol

OpenStudy (anonymous):

then dived 1024 by 2 = 512 so x is 512 ?

OpenStudy (amistre64):

16(16) = 256+4 = 1024...thats right 1024/2 = 512... x = 512. good job ;)

OpenStudy (amistre64):

256*4 ....my keyboard hates me

OpenStudy (anonymous):

THANKS YOu

OpenStudy (amistre64):

youre welcome, but you already had it figured out :)

OpenStudy (anonymous):

ha trust me if i did, i wouldnt be on here

OpenStudy (amistre64):

:) you did great ;)